Question 4

Jack is planning to buy a 4-year bond with semi-annual coupons and a coupon rate of 4.2 percent p.a. The face value is $1,000. Given an annual yield of 10.5 percent, what is the bond’s current price? (to the nearest cent)

Select one:

a. $798.45

b. $802.44

c. $1227.61

d. $670.73
Answer #1

bond’s current price = 802.44

cash flow= 1000*4.2%=42

Year Cash flow PV factor r= 1/(1+r)t @ 10.5% PV of cash flow
0 0 1.00000 0.00
1 42 0.90498 38.01
2 42 0.81898 34.40
3 42 0.74116 31.13
4 42 0.67073 28.17
4 1000 0.67073 670.73
NPV 802.44